### v1

@@ IV-D Parameter Sensitivity and Qualitative Analysis
The Impact of $\alpha$ on Model Performance. As shown in Figure 4, we present the trend of average accuracy across all transfer tasks in the e Office-Home dataset as the parameter $\alpha$ varies. A larger $\alpha$ indicates a greater influence of the cross-domain feature transfer branch’s predictions in the final decision. The results reveal that when $\alpha$ is less than 0.7, the combined performance surpasses that of either branch individually. However, when $\alpha$ exceeds 0.8, the combined performance, while still higher than that of the cross-domain transfer branch, falls below that of the target-specific feature learning branch. This is due to the lower performance of the cross-domain transfer branch compared to the target-specific branch within the e Office-Home dataset. Overall, the fusion of the two branches significantly enhances their performance across most values of $\alpha$.
@@
Figure 5: The t-SNE visualization of R $\to$ P transfer tasks in Office-Home. Different colored dots represent the sample features of the target domain, while black pentagrams and red inverted triangles denote categorical text features. (a) Categorical text features derived using the manually designed prompt ”a photo of a” from CLIP; (b) Categorical text features obtained from the source domain after prompt learning; (c) Two different types of categorical text features obtained from our dual-branch network after unsupervised training, where the black pentagrams indicate the cross-domain feature transfer branch and the red inverted triangles indicate the target-specific feature learning branch.
The Impact of $K$ on Model Performance. To analyze the impact of the number of high-confidence target domain samples within the cross-domain feature transfer branch on model performance, we conducted this experiment across 12 transfer tasks within the e Office-Home dataset. We compared the average accuracy of all transfer tasks in the target domain under varying counts of high-confidence samples per category, specifically 1, 2, 4, 8, and 16 samples. As shown in Table IX, when the value of $K$ increased from 1 to 8, there was a significant improvement in model performance, with an increase of 2.29% when using 8 samples per category compared to just one. However, further increasing the count to 16 did not alter performance significantly. This plateau is attributed to the increased likelihood of introducing erroneous pseudo-labels as the number of samples grows, adversely affecting the model’s effectiveness.
@@
TABLE IX: Average Accuracy for Different Values of $K$ on the OfficeHome Dataset Using RN-50 Backbone.
Dual-Branch Derived Categorical Text Features Visualization Using t-SNE. To qualitatively analyze the variation in the class text features obtained, we conducted this experiment. Specifically, we randomly selected a transfer direction from the e Office-Home dataset and, for visualization convenience, randomly chose 65 categories from the 65 available. We compared the class text features derived using the manually designed prompt ”a photo of a” in CLIP, those obtained from the source domain after prompt-learning training, and those derived from our proposed dual-branch network following unsupervised training. Our method incorporates two types of class text features. As shown in Figure 5, the class text features obtained from the manually designed prompt template and source domain prompt-learning are more concentrated, displaying poor class separability for the target domain samples. In contrast, after enhancement through our cross-domain transfer branch, the class text features from the source domain are dispersed across different sample clusters, and the target-specific feature learning branch also distributes class text features across various class clusters, demonstrating improved inter-class separability.
@@ V Conclusion
In this paper, we address the dual challenges of inaccessible source domains and limited source sample availability in FS-SF-UDA by proposing a Data-Efficient CLIP-Powered Dual-Branch Network(CDBN). This network leverages the synergistic effects of cross-domain feature transfer and target-specific feature learning to enhance model generalization to the target domain while preserving the effective class semantic information learned from a small number of source domain samples. This significantly improves data utilization efficiency for source domain samples and prevents model overfitting to the limited source domain sample distribution. Furthermore, we introduce an unsupervised optimization strategy that maintains the source domain’s classification capabilities while ensuring that the target domain predictions balance high accuracy with category diversity. Employing only unlabeled target domain samples, this strategy enables stable and efficient model training. Ultimately, our approach achieves superior performance on multiple datasets and comparable results to the existing best methods on other datasets. This validates the effectiveness of our proposed dual-branch structure and its unsupervised optimization strategy, further demonstrating that even with limited source domain samples, the category semantic information can still provide crucial assistance for target domain adaptation.

### v2

@@ IV-D Parameter Sensitivity and Qualitative Analysis
The Impact of $\alpha$ on Model Performance. As shown in Figure 4, we present the trend of average accuracy across all transfer tasks in the Office-Home dataset as the parameter $\alpha$ varies. A larger $\alpha$ indicates a greater influence of the cross-domain feature transfer branch’s predictions in the final decision. The results reveal that when $\alpha$ is less than 0.7, the combined performance surpasses that of either branch individually. However, when $\alpha$ exceeds 0.8, the combined performance, while still higher than that of the cross-domain transfer branch, falls below that of the target-specific feature learning branch. This is due to the lower performance of the cross-domain transfer branch compared to the target-specific branch within the Office-Home dataset. Overall, the fusion of the two branches significantly enhances their performance across most values of $\alpha$.
@@
Figure 5: The t-SNE visualization of R $\to$ P transfer tasks in Office-Home. Different colored dots represent the sample features of the target domain, while black pentagrams and magenta inverted triangles denote categorical text features. (a) Categorical text features derived using the manually designed prompt ”a photo of a” from CLIP; (b) Categorical text features obtained from the source domain after prompt learning; (c) Two different types of categorical text features obtained from our dual-branch network after unsupervised training, where the black pentagrams indicate the cross-domain feature transfer branch and the magenta inverted triangles indicate the target-specific feature learning branch.
The Impact of $K$ on Model Performance. To analyze the impact of the number of high-confidence target domain samples within the cross-domain feature transfer branch on model performance, we conducted this experiment across 12 transfer tasks within the Office-Home dataset. We compared the average accuracy of all transfer tasks in the target domain under varying counts of high-confidence samples per category, specifically 1, 2, 4, 8, and 16 samples. As shown in Table IX, when the value of $K$ increased from 1 to 8, there was a significant improvement in model performance, with an increase of 2.29% when using 8 samples per category compared to just one. However, further increasing the count to 16 did not alter performance significantly. This plateau is attributed to the increased likelihood of introducing erroneous pseudo-labels as the number of samples grows, adversely affecting the model’s effectiveness.
@@
TABLE IX: Average Accuracy for Different Values of $K$ on the OfficeHome Dataset Using ResNet-50 Backbone.
Dual-Branch Derived Category Text Features Visualization Using t-SNE. To qualitatively analyze the variation in the class text features obtained, we conducted this experiment. Specifically, we randomly selected a transfer direction from the Office-Home dataset and, for visualization convenience, randomly chose 25 categories from the 65 available. We compared the class text features derived using the manually designed prompt ”a photo of a” in CLIP, those obtained from the source domain after prompt-learning training, and those derived from our proposed dual-branch network following unsupervised training. Our method incorporates two types of class text features. As shown in Figure 5, the class text features obtained from the manually designed prompt template and source domain prompt-learning are more concentrated, displaying poor class separability for the target domain samples. In contrast, after enhancement through our cross-domain transfer branch, the class text features from the source domain are dispersed across different sample clusters, and the target-specific feature learning branch also distributes class text features across various class clusters, demonstrating improved inter-class separability.
@@ V Conclusion
In this paper, we address the dual challenges of inaccessible source domains and limited source sample availability in FS-SF-UDA by proposing a Data-Efficient CLIP-Powered Dual-Branch Network(CDBN). This network leverages the synergistic effects of cross-domain feature transfer and target-specific feature learning to enhance model generalization to the target domain while preserving the effective class semantic information learned from a small number of source domain samples. This significantly improves data utilization efficiency for source domain samples and prevents model overfitting to the limited source domain sample distribution. Furthermore, we introduce an unsupervised optimization strategy that maintains the source domain’s classification capabilities while ensuring that the target domain predictions balance high accuracy with category diversity. Employing only unlabeled target domain samples, this strategy enables stable and efficient model training. Ultimately, our approach achieves best performance on multiple datasets and comparable results to the existing best methods on other datasets. This validates the effectiveness of our proposed dual-branch structure and its unsupervised optimization strategy, further demonstrating that even with limited source domain samples, the category semantic information can still provide crucial assistance for target domain adaptation.